# StableQ: Enhancing Data-Scarce Quantization with Text-to-Image Data

### 07 Dec 2023 v1

@@ Abstract
Though low-bit quantization enables efficient storage and inference of deep neural networks, it often requires the use of training data to maintain resilience against quantization errors. However, training data are frequently subject to privacy or copyright concerns. In this work, we address the challenge of Data-Scarce Quantization, where access to training data is severely limited or non-existent for quantization purposes. Conventional approaches typically rely on inverting dummy images or jointly training generative models to produce synthetic input samples. However, these methods struggle to accurately recreate complex objects in large-scale datasets like ImageNet. To overcome these limitations, we introduce StableQ , a novel method that utilizes an advanced text-to-image diffusion model to generate high-resolution, photo-realistic synthetic data. To verify the quality of the generated data, we implement two robust filtering mechanisms. These mechanisms are designed to select images that closely resemble the intrinsic characteristics of the actual training data. Furthermore, in scenarios where limited training data are available, we use these data to guide the synthetic data generation process by inverting a learnable token embedding in the text encoder. Our extensive experimental results demonstrate that StableQ sets a new benchmark in both zero-shot and few-shot quantization, outperforming existing methods in terms of accuracy and efficiency.
@@ 1 Introduction
A variety of model compression techniques have been developed to enable the deployment of deep learning models on embedded/mobile devices without significant performance loss. One prominent method is neural network quantization [22], which involves converting 32-bit floating-point models into compact low-bit fixed-point formats. This transformation leverages the efficiency of fixed-point computation and the benefits of reduced memory usage. Other notable techniques include pruning [6] and knowledge distillation [28], both of which aim to reduce network size while preserving as much of the original model’s performance as possible.
@@
While these methods effectively accelerate neural networks, they typically require finetuning on the original training dataset to mitigate any accuracy loss in the compressed model. However, accessing the original training data can be challenging due to privacy concerns or Intellectual Property (IP) protection issues [47]. This challenge has led to the emergence of data-free quantization [25, 37, 8] (also known as zero-shot quantization). This approach involves generating synthetic data from a pretrained network and then finetuning the compressed or quantized model using this synthetic data, thereby circumventing the need for original training data.
Figure 1: Examples images generated by StableQ on ImageNet.
@@
Several prior works generate synthetic data by inverting the knowledge from pretrained full-precision models. For instance, [8, 25, 72] propose to optimize the images by aligning them with the training data in terms of activation statistics. [65, 43] apply generative adversarial networks [24] to synthesize the images. However, these approaches often struggle to match the compression performance achievable with real training data. The primary challenge lies in the inherent complexity of reverse mapping from a lower to a higher dimension [39]. For example, in the context of the ImageNet dataset, this involves mapping from a 1000-dimensional space to an image space of $224\times 224\times 3$. Such transformation makes accurately recreating complex objects a hard problem.
@@
In this paper, we introduce StableQ , a novel pipeline utilizing the leading open-source text-to-image model, Stable Diffusion [56], to generate high-quality data for finetuning quantized models. This approach aligns with the growing trend of applying AI Generated Content (AIGC) in deep learning and general vision applications [70, 7, 62]. While direct application of text-to-image synthetic data risks distribution shift from original training data, StableQ addresses this through two innovative filtering mechanisms, ensuring the selection of in-distribution synthetic data. StableQ excels in both zero-shot and few-shot quantization scenarios. In few-shot quantization, where limited training data is available, StableQ employs prompt tuning to align synthetic image generation more closely with the real data. This approach marks a significant advancement in quantization-aware training, particularly in data-scarce environments. Fig. 1 showcases example images generated by StableQ, illustrating its efficacy. We conduct comprehensive experiments to validate the effectiveness of StableQ, including both Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT), on various neural network architectures (e.g. CNNs [26, 58] and ViTs [16]).
@@
    We propose StableQ , the first work to leverage advanced text-to-image synthetic data for Data-Scarce Quantization.
@@
    To reduce the distribution gap between synthetic data and real data, we introduce the energy score filter and the Batch Normalization (BN) sensitivity filter to select in-distribution synthetic data. We also propose to learn the dataset token to bridge the gap between few-shot training data and synthetic data.
@@
    Extensive experiments demonstrate StableQ generates state-of-the-art quality data, achieving the best quantization performance in both PTQ and QAT. For example, our 3-bit QAT-based ResNet-50 achieves 74.0% accuracy on the ImageNet dataset, outperforming the existing method by 48%.
@@ 2 Related Work
Quantization Methods. Quantization approaches to compress pretrained Deep Neural Networks (DNNs) can be broadly divided into PTQ and QAT. PTQ performs calibration on a pretrained DNN after quantizing its weights and activations to low bits to maintain the original accuracy [49, 19, 45]. For example, [3, 20] propose bias correction in the convolutional layers after quantization. [74] splits the outliers into additional channels to reduce the quantization error. Recently, a line of works [48, 38, 64, 32] leverages the weight rounding optimization to reconstruct the activation of the original model. While PTQ mostly adopts ad hoc strategies to prevent accuracy loss from quantization, QAT can significantly improve the accuracy of the quantized model through end-to-end finetuning. The success of QAT is mainly based on the usage of Straight-Through Estimator (STE) [68], enabling gradient-based optimization [31, 10, 18, 76, 71, 66]. QAT also explores optimizing the step size [18], clipping threshold [10], non-linear interval [37], non-uniform quantization levels [71] together with the weights.
Data in Quantization. To quantize a model below 8 bits, data is essential, especially in QAT where end-to-end finetuning occurs. When access to real images is restricted due to privacy and copyright issues, zero-shot quantization [8] proposes to synthesize images in replacement of real data. There are two categories of creating synthetic images: (1) Inverting the images directly via gradient descent [8, 72, 39, 25, 36], and (2) Learning a Generative Adversarial Network (GAN) to continuously generate images [65, 43].
@@
Our method opens a third category in data-scarce quantization, i.e., using text-to-image diffusion models. We also investigate the few-shot quantization that generates synthetic data given limited training data to better finetune the quantized model. To the best of our knowledge, this scenario is being studied for the first time.
@@ 3.1 Quantization
There are several algorithms to determine the step size $s$, depending on the specific quantization regime. For instance, under QAT, we utilize the LSQ [18] to update the step size with gradient descent. The gradient to step size can be computed with STE [68], i.e., $\frac{\partial\lfloor x\rceil}{\partial x}=1$. However, under the PTQ regime, the training data is much less than QAT, hindering the learning of step size. To obtain the desired step size under PTQ, we can use the min-max algorithm or quantization-error minimization [3]. In addition, we also optimize weight rounding during PTQ [48, 38, 64, 36], which can significantly improve the model performance.
3.2 DDPM
The Denoising Diffusion Probabilistic Model (DDPM) [29] is a fundamental component of today’s text-to-image models. The DDPM is a parameterized Markov chain, iteratively denoising the noisy images with finite transition steps. During training, the transition kernels are learnt in the reverse direction to perturbing natural images with noise. The noise is added to the data in each step and estimated as the optimization target.
Forward Process. We briefly introduce the Markov process where Gaussian noise is added to the clean image at each step till it is completely randomized. Given a data distribution ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$, DDPM generates ${\bm{x}}_{T}$ iteratively with $q({\bm{x}}_{T}|{\bm{x}}_{T-1})$, given by
$q({\bm{x}}_{1:T}|{\bm{x}}_{0})=\prod_{t=1}^{T}q({\bm{x}}_{t}|{\bm{x}}_{t-1}),$ (3)
where $q({\bm{x}}_{t}|{\bm{x}}_{t-1}):=\mathcal{N}({\bm{x}}_{t};\sqrt{1-\beta}{\bm{x}}_{t-1},\beta I)$. $T$ is number of diffusion steps and $\beta$ is a hyper-parameter to control the perturbation. $\mathcal{N}$ is the Gaussian noise transition kernel.
Reverse Process. Now we can train the transition kernel with a reverse process. Let $\theta$ denote its parameters, we predict the generated distribution $p_{\theta}({\bm{x}}_{0})$ starting from $p_{\theta}({\bm{x}}_{T})$. The optimization objective is given by
$\min_{\theta}\mathbb{E}_{t\sim\mathcal{U}(1,T),{\bm{x}}_{0}\sim q({\bm{x}}_{0}),\epsilon\sim\mathcal{N}(\bf{0},{\bm{I}})}\lambda(t)||\epsilon-\epsilon_{\theta}({\bm{x}}_{t},t)||_{F}^{2},$ (4)
where $\lambda(t)$ is a positive weighting function [29], $\epsilon$ is a noise vector predicted from ${\bm{x}}_{t}$, and $\epsilon_{\theta}$ represents a deep neural network with parameter $\theta$. In this work, we mainly adopt the Stable Diffusion [60], one of the leading open-sourced text-to-image models. In addition to DDPM, the Stable Diffusion further improves the data quality and reduces the complexity by performing diffusion and denoising in the latent space. It uses a VQ-GAN [17] to encode the image into the latent space and decode the latent image after denoising. Also, Stable Diffusion uses cross-model attention to generate images under various forms of conditions like text and images. In StableQ , we mainly adopt Stable Diffusion v1-5 to synthesize the images.
@@ 4 Methodology
In this section, we introduce our methodology to synthesize the training data with Stable Diffusion. We discuss two scenarios in data-scarce quantization: (1) Zero-Shot Quantization (ZSQ) where only the label space and a pretrained full-precision model are provided, and (2) Few-Shot Quantization (FSQ) where limited training data is provided in addition to the pretrained model and label space. For ZSQ, we consider both PTQ and QAT, while for FSQ, we consider only QAT. We first discuss StableQ ’s data generation method under ZSQ and FSQ and then introduce the quantization techniques used in StableQ.
Figure 2: The overall image synthesis and filtering procedure of StableQ. \⃝raisebox{-0.9pt}{1} For zero-shot synthesis, we first generate synthetic images from the stable diffusion. \⃝raisebox{-0.9pt}{2} Based on the generated synthetic images, we compute BN distance and Energy score, then use them for filtering out low-quality images. \⃝raisebox{-0.9pt}{3} For few-shot synthesis, few-shot real images are used to optimize prompt $\{\mathtt{S}\}$ (Eq. (missing) 10). We then generate the synthetic images from the stable diffusion with the optimized prompt. \⃝raisebox{-0.9pt}{4} We utilize the BN distance to train a simple MLP model for filtering images, called Parameterized Filtering.
4.1 Zero-Shot StableQ
We propose a label prompt method that only uses the label as the prompt to synthesize the images without accessing the pretrained model. For example, for an ImageNet-pretrained model [13], we can synthesize the image based on 1000 class names.
@@ 4.1 Zero-Shot StableQ
where {D} is a template adjective (e.g., “nice, dark, small”) derived from the CLIP ImageNet template [54] [^1] We list the full template in Appendix A. {C} is a randomly chosen class label name. As an example, suppose we choose the class hamster, then a random prompt can be generated as "A photo of a clean hamster". Subsequently, we directly use the text-to-image model like Stable Diffusion to synthesize the data.
In our label prompt method, we generate the images relying solely on the label name of each class and the ability of Stable Diffusion (Fig. 2\⃝raisebox{-0.9pt}{1}). However, this operation does not leverage any prior knowledge in the original training data. As a result, the generated image would not be accurate if the target dataset has a large domain gap with the training data of Stable Diffusion or if the label annotation in the target dataset cannot accurately describe the objects. To generate better-quality data for quantization, we propose model-dependent selection to leverage the prior knowledge embedded in the pretrained full-precision model.
We propose two filtering mechanisms for model-dependent selection. We argue that the synthetic data selection can be treated as an Out-of-Distribution (OOD) detection problem [51]. However, contrary to OOD detection, we try to select in-distribution data among the synthetic data for quantization purposes (Fig. 2\⃝raisebox{-0.9pt}{2}).
Energy Filtering. We use energy filtering, which calculates the energy score [42], given by
$E({\bm{x}},f)=-T\sum_{i=1}^{C}e^{-f_{i}({\bm{x}})/T},$ (6)
where $f_{i}({\bm{x}})$ denotes the $i$-th value of the network’s output logits. $C$ and $T$ denote the number of image classes and temperature, respectively. Liu et al. [42] have shown that the energy score can be used as an OOD detector due to its theoretical connection with the likelihood function. The neural networks trained by cross-entropy loss inherently decrease the energy of the training data, hence, the OOD data will have relatively higher energy.
To select synthetic images that have a similar quantization effect to that of real image data, we pass the synthetic data through the full-precision pretrained model. Then, we calculate the energy scores of all synthetic images and only select those images for quantization that yield energy scores lower than a certain user-defined threshold. We will experiment with the choice of this threshold in Sec. 5.6.
BatchNorm Distribution Filtering. Originally proposed in [67, 8], the channel mean and channel variance distance computed in the Batch Normalization (BN) layers [34] are regarded as a golden metric for optimizing the synthetic data. [15] also shows that the channel mean discrepancy can also be an OOD detector. We use this metric to evaluate the distance between synthetic images and original training images on the full-precision pretrained model. We calculate the BN distance by
@@
where $\mu_{l},\sigma_{l}$ denote the running mean and variance of the layer $l$ activation from original training images, and $\mu^{s}_{l},\sigma^{s}_{l}$ are the current mean and variance of the layer $l$ activation from synthetic images. Naively, for each generated synthetic data, we can evaluate the BN distance $D_{BN}$ and filter out the ones with large $D_{BN}$. However, we find this approach does not bring improvement in the final data quality due to the difference between single data and batched data. Evaluating $D_{BN}$ on a single image will lead to biased estimation as it ignores its interaction with other data. Ioffe [33] also demonstrates that with small batch sizes, the estimates of the batch mean and variance used during training become a less accurate approximation of the mean and variance used for testing.
@@
Patch Similarity Filtering for ViTs. As we discussed in Sec. 2, since ViTs do not have BN layers, we adopt the patch similarity metric in [40] to do additional filtering of the synthetic images after the energy filtering. ViT quantization is then performed using the filtered synthetic data. We refer to this metric in the Appendix.
4.2 Few-shot StableQ
In this section, we describe how to synthesize and select synthetic data given a limited amount of training data (i.e., few-shot quantization). This case could be more common than the completely zero-shot scenario as in practice it is not hard to obtain some permitted training data. We hypothesize that synthesizing a number of images based on this limited dataset can enhance quantization performance.

### 08 Mar 2024 v2

# GenQ: Quantization in Low Data Regimes with Generative Synthetic Data

@@ Abstract
In the realm of deep neural network deployment, low-bit quantization presents a promising avenue for enhancing computational efficiency. However, it often hinges on the availability of training data to mitigate quantization errors, a significant challenge when data availability is scarce or restricted due to privacy or copyright concerns. Addressing this, we introduce GenQ, a novel approach employing an advanced Generative AI model to generate photorealistic, high-resolution synthetic data, overcoming the limitations of traditional methods that struggle to accurately mimic complex objects in extensive datasets like ImageNet. Our methodology is underscored by two robust filtering mechanisms designed to ensure the synthetic data closely aligns with the intrinsic characteristics of the actual training data. In case of limited data availability, the actual data is used to guide the synthetic data generation process, enhancing fidelity through the inversion of learnable token embeddings. Through rigorous experimentation, GenQ establishes new benchmarks in data-free and data-scarce quantization, significantly outperforming existing methods in accuracy and efficiency, thereby setting a new standard for quantization in low data regimes.
@@ 1 Introduction
A variety of model compression techniques have been developed to deploy large deep learning models on embedded/mobile devices without significant accuracy drops. One prominent method is neural network quantization [18], which involves converting 32-bit floating-point models into a compact low-bit fixed-point format. This transformation leverages the efficiency of fixed-point computation and the benefits of reduced memory usage. Other notable techniques include pruning [5, 44] and knowledge distillation [24], both of which aim to reduce network size while preserving as much of the original model’s performance as possible.
@@
While these methods effectively accelerate neural networks, they typically require finetuning on the original training dataset to mitigate any accuracy loss in the compressed model. However, accessing the original training data can be challenging due to privacy concerns or Intellectual Property (IP) protection issues [43]. This challenge has led to the emergence of data-free quantization [21, 33, 7]. This approach involves generating synthetic data from a pretrained network and then finetuning the compressed or quantized model using this synthetic data, thereby circumventing the need for original training data.
Figure 1: Comparison of GenQ with existing methods on ImageNet. (1) Data-Free PTQ, (2) Data-Free QAT, (3) Data Generation Speed. Real* denotes using real ImageNet data in zero-shot quantization.
@@
Several prior works generate synthetic data by inverting the knowledge from pre-trained full-precision models. For instance, [7, 21, 69] propose to optimize the images by aligning them with the training data in terms of activation statistics. [62, 39] apply generative adversarial networks [20] to synthesize the images. However, these approaches often struggle to match the compression performance achievable with real training data. The primary challenge lies in the inherent complexity of reverse mapping from a lower to a higher dimension [35]. For example, in the context of the ImageNet dataset, this involves mapping from a 1000-dimensional space to an image space of $224\times 224\times 3$. Such transformation makes accurate recreation of complex objects a hard problem. Additionally, these methods cost a significant amount of time to generate the images.
@@
In this paper, we introduce GenQ, a new pipeline utilizing the advanced Generative AI model such as Stable Diffusion [56] to generate high-quality data for quantization purposes. This approach aligns with the growing trend of applying AI-generated content (AIGC) in deep learning and vision applications [67, 6, 58, 61]. While direct application of generative AI data risks distribution shifts from original training data, we address this through two filtering mechanisms, ensuring the selection of in-distribution synthetic data. Our method excels in data-free quantization scenarios. More remarkably, in a majority of the model deployment scenarios with limited training data (e.g. one image per class), GenQ is significantly effective via prompt tuning to align generated images more closely with the real data. We conduct comprehensive experiments to validate the effectiveness of GenQ, including both Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT), on various neural network architectures (e.g. CNNs [22, 54] and ViTs [14]). Fig. 1 showcases effectiveness of GenQ on data-free quantization. Meanwhile, compared to other data generation approaches, our method achieves faster generation of synthetic data (up to 15$\times$ faster), increasing both efficiency and effectiveness.
@@
    We propose GenQ, the first work to leverage advanced text-to-image synthetic data for quantization in data-free scenarios. To reduce the distribution gap between synthetic data and real data, we introduce a suite of filtering mechanisms to select in-distribution synthetic data, including energy score filtering and BatchNorm Distribution filtering.
@@
    In case of limited data availability, we also propose to learn the token embedding to guide the synthesis using real data.
@@
    We conduct extensive experiments to show the usefulness of GenQ in generating synthetic data that helps achieve SoTA quantization performance in both PTQ and QAT. For example, our 4-bit QAT-based ResNet-50 achieves 76.10% accuracy on the ImageNet dataset, outperforming the latest existing method [8] by 5.4%.
@@ 2 Related Work
Quantization Methods. Quantization approaches to compress pretrained Deep Neural Networks (DNNs) can be broadly divided into PTQ and QAT. PTQ performs calibration on a pre-trained DNN after quantizing its weights and activations to low bits to maintain the original accuracy [46, 16, 41]. For example, [3, 17] propose bias correction in the convolutional layers after quantization. [71] splits the outliers into additional channels to reduce the quantization error. Recently, a line of works [45, 34, 60, 28] leverage the weight rounding optimization to reconstruct the activation of the original model. While PTQ mostly adopts ad hoc strategies to prevent accuracy loss from quantization, QAT can significantly improve the accuracy of the quantized model through end-to-end finetuning. The success of QAT is mainly based on the usage of Straight-Through Estimator (STE) [65], enabling gradient-based optimization [27, 9, 15, 73, 68, 63]. QAT also explores optimizing the step size [15], clipping threshold [9], non-linear interval [33], non-uniform quantization levels [68] together with the weights.
Data in Quantization. To quantize a model to below 8 bits, data is essential, especially in QAT where end-to-end finetuning occurs. When access to real images is restricted due to privacy and copyright issues, ZeroQ [7] proposes to synthesize images in replacement of real data. There are two categories of creating synthetic images: (1) Inverting the images directly via gradient descent [7, 69, 35, 21, 32], and (2) Learning a Generative Adversarial Network (GAN) to continuously generate images [62, 39].
@@
Our method opens a third category for quantization in low data regime, i.e., using text-to-image diffusion models. Additionally, we investigate how to use limited real data to guide the generation of synthetic data. To the best of our knowledge, this scenario is being studied for the first time.
@@ 3.1 Quantization
Quantization-Aware Training (QAT). We utilize the Learned Step Size Quantization (LSQ) [15] to update the step size $s$ and ${\bm{w}}$ with gradient descent. The gradient to step size can be computed with Straight-Through Estimator (STE) [65], i.e., $\frac{\partial\lfloor x\rceil}{\partial x}=1$. QAT requires a significant amount of data and GPU resources to perform end-to-end fine-tuning.
Post-Training Quantization (PTQ). Under this PTQ regime, the training data is much less than QAT, hindering the learning of step size. To obtain the desired step size $s$ under PTQ, we perform layerwise quantization-error minimization [3]. In addition, we also conduct layer-wise tuning of the weight rounding during PTQ as did in [45, 34, 60, 32], which can significantly improve the model performance with only 1K images.
3.2 Stable Diffusion Data Generation
Stable Diffusion utilizes the Denoising Diffusion Probabilistic Model (DDPM) [25] to implement the training and inference. It contains a text encoder $\tau_{\theta}(\cdot)$, image encoder $\mathcal{E}(\cdot)$ and image decoder $\mathcal{D}(\cdot)$, and finally a denoising U-Net $\epsilon_{\theta}(\cdot)$.
Inference Process. During inference, a random noise image latent is sampled from Gaussian distribution ${\bm{z}}\sim\mathcal{N}(\mathbf{0,1})$. Suppose the text prompt is ${\bm{s}}$, the denoising U-Net fuses the text embedding $\tau_{\theta}({\bm{s}})$ and the image and visual embedding through cross-attention layers and denoises the image latent gradually. After $T$ timesteps of denoising, the image latent is decoded by $\mathcal{D}(\cdot)$ to generate the high-resolution image.
Training process. Training the denoising U-Net essentially involves encoding the images first ($\hat{{\bm{x}}}_{0}=\mathcal{E}({\bm{x}})$), and then diffusing the image through a Markov chain process $q(\hat{\bm{x}}_{t}|\hat{\bm{x}}_{t-1})=\mathcal{N}(\hat{\bm{x}}_{t};\sqrt{1-\beta}\hat{\bm{x}}_{t-1},\beta I)$. Note that $t\in[1,T]$ is the number of diffusion timesteps and $\beta$ is a hyper-parameter to control the perturbation. When $t=T$, then $\hat{{\bm{x}}}_{t}={\bm{z}}$. Now the denoising U-Net can be trained by a reverse process by matching the output and the noises added before:
$\min_{\theta}\mathbb{E}_{t\sim\mathcal{U}(1,T),\epsilon\sim\mathcal{N}(\bf{0},{\bm{I}})}\lambda(t)||\epsilon-\epsilon_{\theta}(\hat{\bm{x}}_{t},t)||_{F}^{2},$ (3)
where $\lambda(t)$ is a positive weighting function [25], $\epsilon$ is a noise vector predicted from ${\bm{x}}_{t}$. We use Stable Diffusion v1-5 [56] in our pipeline. We refer to [52] for more details about its principle.
@@ 4 Methodology
In this section, we introduce our methodology to synthesize the training data with Stable Diffusion. We discuss two quantization scenarios in low data regimes: (1) Data-Free Quantization (DFQ) where only the label space and a pre-trained full-precision model are provided, and (2) Data-Scarce Quantization (DSQ) where limited training data is provided in addition to the pre-trained model and label space. We first discuss GenQ’s data generation method under DFQ and DSQ and then introduce the quantization techniques we used.
Figure 2: The overall image synthesis and filtering procedure of GenQ. \⃝raisebox{-0.9pt}{1} For DFQ synthesis, we directly use the label as the text prompt, and \⃝raisebox{-0.9pt}{2} we use several metrics to filter our out-of-distribution synthetic images. \⃝raisebox{-0.9pt}{3} For data-scarce synthesis, real images are used to optimize prompt $\{\mathtt{S}\}$ (Eq. (missing) 11). We then generate the synthetic images with the optimized prompt.
4.1 Data-Free GenQ
We propose a label prompt method that only uses the label as the prompt to synthesize the images without accessing the pre-trained model. For example, for an ImageNet pre-trained model [12], we can synthesize the images based on 1000 class names.
@@ 4.1 Data-Free GenQ
where {D} is a template adjective (e.g., “nice, dark, small") derived from the CLIP ImageNet template [51] [^1] We list the full template in Appendix A. {C} is a randomly chosen class label name. As an example, suppose we choose the class hamster, then a random prompt can be generated as "A photo of a small hamster". Subsequently, we directly use the text-to-image model like Stable Diffusion [52] to synthesize the data.
In our label prompt method, we generate the images relying solely on the label name of each class and the ability of Stable Diffusion (Fig. 2\⃝raisebox{-0.9pt}{1}). However, this operation does not leverage any prior knowledge in the original training data, which might cause a distribution shift from original training data to synthetic training data. To generate better-quality data for quantization, we propose model-dependent selection to leverage the prior knowledge embedded in the pre-trained full-precision model.
We propose a set of filtering mechanisms for model-dependent selection. We argue that the synthetic data selection can be treated as an Out-of-Distribution (OOD) detection problem [48]. However, contrary to OOD detection, we intend to select in-distribution data from the generated synthetic data for quantization purposes (Fig. 2\⃝raisebox{-0.9pt}{2}).
Energy Filtering. We use energy filtering to calculate the energy score [38] as,
$E({\bm{x}},f)=-\alpha\sum_{i=1}^{C}e^{-f_{i}({\bm{x}})/\alpha},$ (5)
where $f_{i}({\bm{x}})$ denotes the $i^{th}$ value of the network’s output logits. $C$ and $\alpha$ denote the number of image classes and temperature, respectively. [38] has shown that the energy score is an OOD detector due to its theoretical connection with the likelihood function. The neural networks trained by cross-entropy loss inherently decrease the energy of the training data, hence, the OOD data will have relatively higher energy.
To select synthetic images that have a similar quantization effect to that of real image data, we pass the synthetic data through the full-precision pre-trained model. Then, we calculate the energy scores of all synthetic images and only select those images for quantization that yield energy scores lower than a certain user-defined threshold. We will experiment with the choice of this threshold in Sec. 5.4.
BatchNorm Distribution Filtering. Originally proposed in [64, 7], the channel mean and channel variance distance computed in the BN layers [30] are regarded as a standard metric for optimizing the synthetic data in Convolutional Neural Networks (CNNs). We use this metric to evaluate the distance between synthetic images and original training images on the full-precision pre-trained model. We calculate the BN distance by
@@
where $\mu_{l},\sigma_{l}$ denote the running mean and variance of the layer $l$ activation from original training images, and $\mu^{s}_{l},\sigma^{s}_{l}$ are the current mean and variance of the layer $l$ activation from synthetic images. Naively, for each generated synthetic data, we can evaluate the BN distance $D_{BN}$ and filter out the ones with large $D_{BN}$. However, we find this approach does not bring improvement in the final data quality due to the difference between single data and batched data. Evaluating $D_{BN}$ on a single image will lead to biased estimation as it ignores its interaction with other data. Ioffe [29] also demonstrates that with small batch sizes, the estimation of the batch mean and variance used during training become a less accurate approximation of the mean and variance used for testing.
@@
Patch Similarity Filtering for ViTs. As we discussed in Sec. 2, since ViTs do not have BN layers, we adopt the patch similarity metric in [36] to do additional filtering of the synthetic images after the energy filtering. Formally, given the output features tensor ${\bm{o}}$, we first calculate the cosine similarity matrices between any two sub-tensors in the patch dimension, given by
[…]
where $\hat{f}_{h}(x)$ is the continuous probability density function of $\mathbf{\Gamma}$, which can be obtained by the kernel density estimation [36]. A low entropy value indicates that the patch similarity distribution is less diverse. Hence, we can select the synthetic data that has the highest diversity in patch similarity.
Filtering Pipeline. The overall filtering pipeline is a two-stage procedure. First, regardless of the type of pre-trained model, we apply energy score filtering. Second, based on the category of the model (CNNs or ViTs), we either apply BatchNorm Distribution Filtering or Patch Similarity Filtering. The selected synthetic data is then used for quantization.
4.2 Data-Scarce GenQ
In this section, we describe how to synthesize and select synthetic data given a limited amount of training data (i.e., data-scarce quantization). This case could be more common than the DFQ scenario as in practice it is not hard to obtain some permitted training data. We hypothesize that synthesizing a number of images based on this limited dataset can enhance quantization performance.